Question 1

1-calculate the market equilibrium if the equation for a market demand curve is Qd=100-0,5P and the equestion for a market supply curve is Qs= -20 +P . (6 marks)
1
Expert's answer
2020-03-17T09:26:59-0400

In equilibrium, Qd = Qs

- 20+P=100-0,5

1,5P=120

 P = 80

Therefore equilibrium quantity , Q = Qd = Qs = 100 - 0,5х80 = 60 unit

 Equilibrium price is 80 unit and quantity is 60 unit.